Terephtgalic acid is an important chemical used in the manufacture of polysters and plasticizers. It contains only C,H, and O. Combustion of 19.81 mg of terephthalic acid produced 41.68 mg CO2 and 6.45 mg H2O. The molar mass of terephthalic acid is 166 g. Calculate the empirical and the molecular formula of terephthalic acid.
1
Expert's answer
2020-11-22T11:13:47-0500

Chemical reaction:

"C_xH_yO_z+O_2\iff CO_2+H_2O"


m"(C_xH_yO_z)"= 19.81 mg

m"(CO_2)" = 41.68 mg

m"(H_2O)" = 6.45 mg


mass of carbon:

"44-41.68m g"

"12- x=11.3672mg"


mass of hydrogen:

"18 - 6.45mg"

"2-x=0.7167g"


mass of oxygen:

m(oxygen) = m(terephtgalic acid) - m(carbon) - m(hydrgen) = 7.7261mg


"(C_xH_yO_z) = \\dfrac{11.3672}{12}\t= \\dfrac{0.7167}{1}\t= \\dfrac{7.7261}{16}\t=" "0.9472647:0.7167:0.48288125"


need to devide each number by the smallest one:


1.96169286:1.48421584:1

"n(C_xH_yO_z)=166g" = 166g

"41n=166g"

"n=4"


need to multiply above ratio by 4, then we get molecular formula of terephtgalic acid:


8:6:4

"C_8H_6O_4"


Since all three coefficients have the common factor 2, divide the molecular formula by 2 to find the empirical formula:

"C_4H_3O_2"